The starter motor of a car engine draws a current of 180A from the battery. The copper wire to the motor is 4.20mm in diameter and 1.2 m long. The starter motor runs for 0.830s until the car engine starts.

A) How much charge passes through the starter motor?

B) How far does an electron travel along the wire while the starter motor is on?


Q2:What is the resistance of
A) A 1.90

m long copper wire that is 0.100mm in diameter?

B) A 50.0

cm long piece of carbon with a 2.00mm
Answer #1

A) Q = i*t = ne

n = i*t/e= 180*0.83/(1.6*10^-19) =933.75e+18


B) d = v*t = I*t/neA = (180*0.83)/(933.75e+18*1.6*10^-19*3.14*(2.1*10^-3)^2) = 72.2157e+3 m